Exercise 17.96 Calculate the molar solubility of Cux (Ksp 1.27 x 10 n each of the following Correct Part B 0.28 M CuCl2 Express your answer using two significant figures. Submit My Answers Give Up Part C 0.23 M Na2x Express your answer using two significant figures. Submit My Answers Give Up Provide Feedback Continue
Answer #1

CuX will have the next dissolutoin

CuX <-> Cu+2 + X-2

Ksp = [Cu+2][X-2]

if

CuCl2 is added, then

[Cu+2] = [CuCl2] = 0.28 M

so

Ksp = [Cu+2][X-2]

1.27*10^-36 = (0.28)(S)

S = (1.27*10^-36) / 0.28 = 4.5357*10^-36 M

C)

for

Na2X --< 2Na+ and X-2

so

[X-2] = 0.23 M

Ksp = [Cu+2][X-2]

1.27*10^-36 = (cu+2)(0.23 )

Cu+2 = (1.27*10^-36) / 0.23

Cu+2 = 5.5217*10^-36 M